# MEGC2025: Micro-Expression Grand Challenge on Spot Then Recognize and Visual Question Answering

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27-October 31, 2025; Dublin, Irelandisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Computer visionccs: Applied computing Psychology
@@
Table 1. The test set results of the ME-STR task using MEAN trained on CAS(ME)2. STRS refers to the spot-then-recognize score.
1. Introduction
Over the years, we have organized a series of ME grand challenges (MEGCs) to explore emerging directions and advance the state of the art in ME research. To date, seven MEGCs have been held: FG’18 (Yap et al., 2018), focusing on ME recognition; FG’19 (See et al., 2019), addressing both ME spotting and recognition; FG’20 (Jingting et al., 2020), centered on ME spotting; MM’21 (Li et al., 2021), introducing ME generation and continuing with spotting; MM’22 (Li et al., 2022c), continuing with spotting and generation; MM’23 (Davison et al., 2023), with a focus on spotting; and MM’24 (See et al., 2024), which introduced the spot-then-recognize paradigm and cross-cultural spotting.
Task 1: ME-STR. Since the surge of interest in ME research over the past decade, the community has primarily treated spotting and recognition as separate tasks. However, such separation is often impractical in real-world settings. ME recognition alone assumes that the ME sequence has already been spotted, which is rarely the case in the real world. Conversely, ME spotting without recognition fails to provide a meaningful interpretation of the emotional content. The spot-then-recognize (STR) paradigm offers a more realistic setting by first identifying the occurrence of a ME (spotting), and then classifying it into an ME category (recognition). In this task, only correctly spotted samples are passed on for emotion classification. This task is a continuation from MEGC 2024 (See et al., 2024).
@@ 1. Introduction
Table 2. The test set results of the ME-VQA task using Qwen2.5VL-3B trained on the curated ME-VQA dataset. The coarse emotion class refers to positive, negative, and surprise. The fine-grained emotion class refers to happiness, surprise, fear, disgust, anger, and sadness. OAO refers to the input with onset, apex, and offset frames. OF referees to the optical flow input computed between the onset and apex frames. ZS refers to zero-shot performance, while FT refers to fine-tuned performance.
2. Task 1: Spot-Then-Recognize
2.1. Task Formulation
2.2. Dataset
For the test set of ME-STR, it contains 30 long videos, including 10 long videos from the SAMM Challenge dataset (Yap et al., 2022; Davison et al., 2018) and 20 clips cropped from different videos in CAS(ME)3 (unreleased before). The frame rate for SAMM Challenge dataset is 200fps, and the frame rate for CAS(ME)3 is 30 fps. The participants should test on this unseen dataset.
2.3. Method
We consider the micro-expression analysis network (MEAN) (Liong et al., 2023) as our baseline method. MEAN is a unified neural network architecture designed to perform both ME spotting and recognition in an end-to-end manner. It is structured with shared layers followed by two distinct branches: one for spotting and one for recognition. The spotting branch functions as a regression branch, outputting a confidence score for each frame to indicate the likelihood of an ME interval. Once the spotting branch identifies candidate intervals of MEs, the recognition branch predicts the emotion class for each interval.
2.4. Evaluation Metrics
[…]
The details of these metrics can be found in (Liong et al., 2023).
2.5. Experiments
Table 1 presents the quantitative results of the STR task using the MEAN network training on CAS(ME)2 and testing on the MEGC 2025 test set of ME-STR. The spotting results indicate the substantial challenge of localizing MEs in long video sequences, as evidenced by the high numbers of FPs and FNs, and the low numbers of precision, recall, and F1 in both datasets. The recognition results, evaluated on the intervals detected by the spotting module, demonstrate a slightly higher performance on certain metrics, but are ultimately constrained by the quality of the detected intervals. The overall STRS score, defined as the product of the spotting and recognition F1-scores, remains low for both datasets (0.0062 for SAMM and 0.0086 for CAS(ME)3), primarily due to the limited recall and precision achieved in the spotting stage. This outcome highlights that the bottleneck in ME analysis systems lies predominantly in the spotting task; improvements in the temporal localization of ME are likely to yield substantial gains in end-to-end analysis performance.
3. Task 2: Visual Question Answering
3.1. Task Formulation
@@ 3.2. Dataset
The ME-VQA test set contains 24 ME clips, including 7 clips from the SAMM Challenge dataset (Yap et al., 2022; Davison et al., 2018) and 17 clips cropped from different videos in CAS(ME)3 (unreleased before). The frame rate for SAMM is 200 fps and the frame rate for CAS(ME)3 is 30 fps. The participants should test on this unseen dataset.
3.3. Method
For the ME-VQA task, we employ the Qwen2.5VL-3B model (Bai et al., 2025) as our baseline method. Qwen2.5VL-3B is a recent LVLM that is based on the Qwen 2.5 architecture and incorporates a vision encoder, a language model backbone, and a cross-modal fusion module. The model is pre-trained on large-scale image-text datasets, enabling it to perform a wide range of VQA tasks with strong zero-shot and few-shot generalization capabilities.
@@ 3.4. Evaluation Metrics
3.5. Experiments

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3762065isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Computer visionccs: Applied computing Psychology
1. Introduction
Figure 1. An overview of Micro-Expression Grand Challenges (MEGCs).
Over the years, we have organised a series of ME grand challenges (MEGCs) to explore emerging directions and advance the state-of-the-art in ME research. To date, seven MEGCs have been held: FG’18 (Yap et al., 2018), focusing on ME recognition; FG’19 (See et al., 2019), addressing both ME spotting and recognition; FG’20 (Jingting et al., 2020), centered on ME spotting; MM’21 (Li et al., 2021), introducing ME generation and continuing with spotting; MM’22 (Li et al., 2022c), continuing with spotting and generation; MM’23 (Davison et al., 2023), with a focus on spotting; and MM’24 (See et al., 2024), which introduced the spot-then-recognize paradigm and cross-cultural spotting. An overview of all the MEGCs can be found in Fig. 1.
Task 1: ME-STR. Since the surge of interest in ME research over the past decade, the community has primarily treated spotting and recognition as separate tasks. However, such separation is often impractical in real-world settings. ME recognition alone assumes that the ME sequence has already been spotted, which is rarely the case in the real world. Conversely, ME spotting without recognition fails to provide a meaningful interpretation of the emotional content. The spot-then-recognize (STR) paradigm offers a more realistic setting by first identifying the occurrence of an ME (spotting), and then classifying it into an ME category (recognition). In this task, only correctly spotted samples are passed on for emotion classification. This task is a continuation from MEGC 2024 (See et al., 2024).
Table 1. Baseline results of the ME-STR task. The test set results of the ME-STR task using MEAN (the baseline method) trained on CAS(ME)2. STRS refers to the spot-then-recognize score.
@@ 1. Introduction
Table 2. Challenge evaluation of the ME-STR task. The test set results of the ME-STR task from the top-performing teams. STRS refers to the spot-then-recognize score.
Participant SAMM (F1) CAS(ME)3 (F1) STRS
Spotting Recognition Spotting Recognition SAMM CAS(ME)3 Overall
Guo et al. (Guo et al., 2025)
0.086 0.667 0.099 0.645 0.057 0.064 0.09
ustc-iat 0.118 0.471 0.067 0.645 0.055 0.043 0.06
gormanv 0.067 0.622 0.061 0.278 0.041 0.017 0.047
2. Task 1: Spot-Then-Recognize
2.1. Task Formulation
2.2. Dataset
For the ME-STR test set, it contains 30 long videos, including 10 long videos from the SAMM Challenge dataset (Yap et al., 2022; Davison et al., 2018) and 20 clips cropped from different videos in CAS(ME)3 (unreleased before). The frame rate for the SAMM Challenge dataset is 200fps, and the frame rate for CAS(ME)3 is 30 fps. Participants should test on this unseen dataset.
2.3. Evaluation Metrics
Several performance metrics are used in this work to ensure fair comparisons and to provide a comprehensive evaluation of the ME spotting and recognition tasks. For ME spotting, true positive (TP), false positive (FP), false negative (FN), precision, recall, and F1 are used. For ME recognition, TP, FP, FN, precision, recall, F1, unweighted F1 (UF1), and unweighted (UAR) are used. Finally, to comprehensively evaluate the performance of the complete ME analysis system, the spot-then-recognize score (STRS) is employed as the product of the F1-scores of spotting ($\mathrm{F1\text{-}score}_{s}$) and analysis ($\mathrm{F1\text{-}score}_{a}$):
The details of these metrics can be found in (Liong et al., 2023).
For our baseline method, we include detailed evaluation metrics. To simplify the challenge evaluation, the methods from the challenge participants will be mainly evaluated based on $\mathrm{F1\text{-}scores}$ and STRS.
2.4. Base line Method
We used the micro-expression analysis network (MEAN) (Liong et al., 2023) as our baseline method. MEAN is a unified neural network architecture designed to perform both ME spotting and recognition in an end-to-end manner. It is structured with shared layers followed by two distinct branches: one for spotting and one for recognition. The spotting branch functions as a regression branch, outputting a confidence score for each frame to indicate the likelihood of an ME interval. Once the spotting branch identifies candidate intervals of MEs, the recognition branch predicts the emotion class for each interval.
2.5. Baseline Results
Table 1 presents the quantitative results of the STR task using the MEAN network training on CAS(ME)2 and testing on the MEGC 2025 ME-STR test set. The spotting results indicate the substantial challenge of localising MEs in long video sequences, as evidenced by the high number of FPs and FNs, and the low numbers of precision, recall, and F1 in both datasets. The recognition results, evaluated on the intervals detected by the spotting module, demonstrate slightly higher performance on certain metrics, but are ultimately constrained by the quality of the detected intervals. The overall STRS score, defined as the product of the spotting and recognition F1-scores, remains low for both datasets (0.0062 for SAMM and 0.0086 for CAS(ME)3), primarily due to the limited recall and precision achieved in the spotting stage. This outcome highlights that the bottleneck in ME analysis systems lies predominantly in the spotting task; improvements in the temporal localisation of ME are likely to yield substantial gains in end-to-end analysis performance.
2.6. Challenge Evaluation
This year, we had 47 registered participants, with 8 participants submitting to the leaderboard for the ME-STR task. The test set results of the ME-STR task from the top-3 teams are shown in Table 2.
Guo et al. (guoguo01) (Guo et al., 2025) introduces a prior-guided video-level regression framework for the ME-STR task, which achieved 1st place on the ME-STR task. It aims to address issues with manually selecting predefined windows for classification and analysis of MEs. The framework integrates scalable interval selection and synergistic optimisation to improve both spotting and recognition of MEs in long video sequences. It also introduces an optimisation framework that works to create synergy between spotting and recognition. This method achieved an overall STRS score of 0.09, and also achieved the best results in almost every other metric for the individual datasets.
The participant ustc-iat introduces frequency-gated spatiotemporal learning (FGSL), a novel framework for ME-STR, achieving the second place in terms of overall STRS. FGSL contains 3 core contributions: a spatiotemporal information gating adapter (IGA) to enhance dependency modelling and attempt to prevent overfitting, a pyramid-based task decoupling to separate localisation and classification, and finally a frequency-based information filtering module (FIFM) for selective feature retention in different scales. Although this method achieved 2nd place overall, it achieved the highest F1 score for spotting on the SAMM dataset. In addition, the method tied top for the analysis F1-score on CAS(ME)3.
The participant gormanv introduces a unified framework for ME analysis using a spatiotemporal multimodal large language model (STM-LLM). It achieved 3rd place in the ME-STR task. It aims to overcome the traditional separation of spotting and recognition tasks by leveraging MLLMs to jointly analyse emotional changes between two frames in video sequences. The method achieved an overall STRS score of 0.047. It also showed promise with the second-highest F1-score on the SAMM analysis.
3. Task 2: Visual Question Answering
3.1. Task Formulation
Table 3. Baseline results of the ME-VQA task. The test set results of the ME-VQA task using Qwen2.5VL-3B (the baseline method) trained on the curated ME-VQA dataset. The coarse emotion class refers to positive, negative, and surprise. The fine-grained emotion class refers to happiness, surprise, fear, disgust, anger, and sadness. OAO refers to the input with onset, apex, and offset frames. OF refers to the optical flow input computed between the onset and apex frames. ZS refers to zero-shot performance, while FT refers to fine-tuned performance.
@@ 3.1. Task Formulation
Table 4. Challenge evaluations of the ME-VQA task. The test set results of the ME-VQA task from the top-performing teams. The coarse emotion class refers to positive, negative, and surprise. The fine-grained emotion class refers to happiness, surprise, fear, disgust, anger, and sadness.
Participant SAMM CAS(ME)3 Overall
Coarse Fine-Grained Coarse Fine-Grained Coarse Fine-Grained
UF1 UAR UF1 UAR BLEU ROUGE UF1 UAR UF1 UAR BLEU ROUGE UF1 UAR UF1 UAR BLEU ROUGE Avg
Wang et al. (Wang et al., 2025) (lianxin-tech) 0.778 0.833 0.427 0.500 0.543 0.625 0.717 0.717 0.294 0.389 0.643 0.605 0.733 0.722 0.368 0.408 0.615 0.607 0.575
Zhu et al. (Zhu et al., 2025) (ustc-iat) 0.750 0.750 0.660 0.700 0.609 0.637 0.561 0.626 0.217 0.278 0.589 0.458 0.594 0.650 0.316 0.375 0.595 0.509 0.506
IIM, HFIPS, CAS 0.675 0.750 0.143 0.143 0.427 0.450 0.438 0.444 0.338 0.333 0.364 0.505 0.560 0.528 0.281 0.283 0.396 0.489 0.423
@@ 3.2. Dataset
The ME-VQA test set contains 24 ME clips, including 7 clips from the SAMM Challenge dataset (Yap et al., 2022; Davison et al., 2018) and 17 clips cropped from different videos in CAS(ME)3 (unreleased before). The frame rate for SAMM is 200 fps, and the frame rate for CAS(ME)3 is 30 fps. The participants should test on this unseen dataset.
@@ 3.3. Evaluation Metrics
The final ranking of the challenge participants will be based on the average (Avg) of all the overall metrics.
3.4. Baseline Method
For the ME-VQA task, we employ the Qwen2.5VL-3B model (Bai et al., 2025) as our baseline method. Qwen2.5VL-3B is a recent LVLM that is based on the Qwen 2.5 architecture and incorporates a vision encoder, a language model backbone, and a cross-modal fusion module. The model is pre-trained on large-scale image-text datasets, enabling it to perform a wide range of VQA tasks with strong zero-shot and few-shot generalisation capabilities.
3.5. Baseline Results
@@
3.6. Challenge Evaluation
The ME-VQA task attracted 28 participants, resulting in 10 teams being submitted to the leaderboard with 259 valid submissions. Among the 10 teams, 9 reported results that exceeded the baseline. The test set results of the ME-VQA task from the top-3 teams are shown in Table 4.
Wang et al. (lianxin-tech) (Wang et al., 2025) proposed Emotion-Qwen-VL and won first place in the ME-VQA task in the MEGC 2025 challenge. Emotion-Qwen-VL is a fully fine-tuned MLLM designed for the ME-VQA. Built on Qwen2.5-VL-7B, the model is optimised to understand subtle facial dynamics and generate interpretable natural language responses. The authors also introduced a rule-based AU-to-emotion mapping framework during the inference to provide a semantic bridge between low-level facial muscle activations and high-level expression labels. In addition, the authors found that training on CASME II and SAMM datasets led to better performance on the SAMM test dataset, while training on CASME II and SMIC datasets led to better performance on the CAS(ME)3 test dataset.
Zhu et al. (ustc-iat) (Zhu et al., 2025) proposed HierMEQA and won second place in the ME-VQA task of the MEGC 2025 challenge. HierMEQA is a novel hierarchical framework designed for the emerging task of ME-VQA. The framework employs a multi-stage hierarchical reasoning process from coarse-grained emotion analysis, fine-grained emotion analysis, to localised action unit detection. In addition, the authors used a hybrid sampling strategy by sampling onset, apex, offset, and some random frames at the training time, while uniform sampling was used at the test time. Instead of full parameter fine-tuning, this method uses parameter-efficient fine-tuning with LoRA.
Participant IIM, HFIPS, CAS proposed a temporal information enhanced VLM (TIE) and won third place in the ME-VQA task in the MEGC 2025 challenge. The proposed framework is built upon Qwen2.5-VL-7B-Instruct and integrates raw video frames and optical flow maps to capture both static and dynamic facial cues. It also used structured prompts to guide the model to focus on specific regions during inference. Full-parameter supervised fine-tuning is used to adapt the model for ME-VQA.
4. Conclusions
In this year’s MEGC challenge, we successfully organised two tasks: ME-STR and ME-VQA. A variety of teams participated, proposing diverse approaches that integrated MLLMs, advanced visual features, and reasoning strategies tailored for ME analysis. Participants exceeded the baseline results by significant margins in both tasks, demonstrating notable improvements in ME-STR and ME-VQA.
For the ME-STR task, the overall STRS remains relatively low, with the best-performing team achieving a score of 0.09. This highlights a substantial need for further exploration and methodological innovation in this area. In contrast, for the ME-VQA task, the best-performing team achieved an average score of 0.575, indicating strong performance. However, the ME-VQA test set for this year’s challenge was relatively small, suggesting that larger and more diverse test datasets will be necessary to better evaluate and validate the generalisation capabilities of ME-VQA methods.
In line with the requirements of ACM Multimedia, only three papers were accepted for publication, one from the STR task and two from the ME-VQA task, based on the performance of the leaderboard, the methodological quality, and the rigor of the submitted papers.
Acknowledgements.
This research is supported in part by the National Natural Science Foundation of China (62476269, 62276252), the Youth Innovation Promotion Association CAS, and the Ministry of Higher Education (MOHE) Malaysia under the Fundamental Research Grant Scheme (FRGS) (FRGS/1/2022/ICT02/HWUM/02/1).